OpenStudy (anonymous):

Y varies directly as x. If y=5/17 when x=10, find when x=5

OpenStudy (anonymous):

5/17 is to 10 as what is to 5? Set this up as a pair of equivalent ratios: \[\frac{ \frac{ 5 }{ 17 } }{ 10 } = \frac{ y }{ 5 }\] Do you know how to solve that? (Hint: cross multiply)